# Efficient Adaptive Experimentation with Non-Compliance

### v1

@@ 6 Theoretical Guarantees
This section provides theoretical guarantees for the AMRIV estimator. We establish its asymptotic normality, characterize its convergence rates, and demonstrate its multiply-robust consistency. Furthermore, in Appendix B, we consider the sequential inference setting and derive asymptotically-valid, time-uniform confidence sequences for the AMRIV estimator.
@@ 6.1 Efficiency and Asymptotic Normality of the AMRIV Estimator
Suppose Assumptions 1, 2 and 3 hold and there exists a non-adaptive policy $\pi(X)\in[\epsilon,1-\epsilon]$ for some $\epsilon>0$ such that the nuisances estimates $\widehat{\eta}_{t}$ and the adaptive assignment policy $\pi_{t}(X\mid\mathcal{H}_{t-1})$ are $L_{2}$-consistent relative to the truncation schedule, i.e. $k_{t}\|\widehat{f}_{t-1}-f\|_{2}=o_{p}(1)$ and $k_{t}\|\pi_{t}-\pi\|_{2}=o_{p}(1)$ for $f\in\{\mu^{Y}(0,\cdot),\mu^{A}(0,\cdot),\delta(\cdot),\delta^{A}(\cdot)\}$. Furthemore, assume $\|\widehat{\delta}_{t-1}-\delta\|_{2}\|\widehat{\delta}^{A}_{t-1}-\delta^{A}\|_{2}=o_{p}(t^{-1/2})$. Then, the AMRIV estimator is asymptotically normal:
[…]
@@
The key insight behind Theorem 3 is that the AMRIV estimator admits the following near-martingale decomposition: $\sqrt{T}(\widehat{\tau}^{\text{AMRIV}}_{T}-\tau)=\sqrt{T}\left(\frac{1}{T}\sum_{t=1}^{T}z_{t}-\tau\right)+\sqrt{T}\left(\frac{1}{T}\sum_{t=1}^{T}m_{t}\right),$ where $z_{t}=\phi(X_{t},Z_{t},A_{t},Y_{t};\pi_{t},\eta)$ is a martingale difference sequence (MDS), and $m_{t}=\phi(X_{t},Z_{t},A_{t},Y_{t};\pi_{t},\widehat{\eta}_{t})-\phi(X_{t},Z_{t},A_{t},Y_{t};\pi_{t},\eta)$ is an asymptotically vanishing term that captures the impact of estimating the nuisance functions. The first term satisfies a central limit theorem for MDS under standard Lindeberg-type conditions [49], while the second is controlled by the $L_{2}$-consistency of the nuisance estimates. We formalize this in Appendix D. Importantly, this result holds under mild assumptions: we only require $L_{2}$ convergence (no pointwise convergence or Donsker conditions Bickel et al. [7]), bounded outcomes and nuisance estimates, and $L_{2}$-consistency of the nuisance components w.r.t. the truncation schedule. This allows AMRIV to accommodate flexible, data-dependent policies and sequential nuisance estimation.
The truncation schedule $k_{t}$ plays a central role by ensuring positivity of $\pi_{t}(X)$—crucial for variance control—while still allowing $\pi_{t}$ to approach an optimal policy $\pi^{*}$ as $k_{t}\to\infty$. For this to hold, we must ensure $\lim_{t\to\infty}k_{t}>\sup_{X}\frac{1}{\pi^{*}(X)}$, so the truncation threshold does not constrain the optimal allocation in the limit and semiparametric efficiency can be achieved (see last line in Theorem 3). This mirrors tradeoffs in efficient ATE estimation [13], where adaptive truncation stabilizes estimation without distorting the estimator asymptotically. In practice, when the plug-in policy $\tilde{\pi}_{t}$ is uniformly bounded away from 00 and $1$, truncation becomes unnecessary: setting $k_{t}=1/\min_{X}\{\tilde{\pi}_{t}(X),1-\tilde{\pi}_{t}(X)\}$ ensures $\pi_{t}=\tilde{\pi}_{t}$ for all $t$.
@@ 6.2 Consistency Guarantees under Partial Nuisance Misspecification
Suppose Assumptions 1, 2 and 3 hold, and that there exists a non-adaptive policy $\pi(X)\in[\epsilon,1-\epsilon]$ for some $\epsilon>0$ such that the adaptive policies $\pi_{t}$ satisfy $k_{t}\|\pi_{t}-\pi\|_{2}=o_{p}(1)$. Let $\widetilde{\eta}=\{\widetilde{\mu}^{Y}(0,\cdot),\widetilde{\mu}^{A}(0,\cdot),\widetilde{\delta}(\cdot),\widetilde{\delta}^{A}(\cdot)\}$ denote a possibly misspecified limit of the nuisance functions, and suppose that $k_{t}\|\widehat{f}_{t-1}-\widetilde{f}\|_{2}=o_{p}(1)$ for each $\widetilde{f}\in\widetilde{\eta}$. Then the AMRIV estimator satisfies
@@
We provide a proof of Theorem 4 and Corollary 5 in Appendix E. An immediate consequence of Theorem 4 is that if both $\widehat{\delta}(X)$ and $\widehat{\delta}^{A}(X)$ converge at rate $o_{p}(T^{-1/4})$, AMRIV achieves the parametric $O_{p}(T^{-1/2})$ rate. This is usually attainable under mild regularity conditions, even with flexible nonparametric models. Furthermore, Corollary 5 shows that AMRIV inherits the multiply robust property from its static counterpart [44, 18]. In the static setting, the MRIV converges if either (i) $\widehat{\mu}^{Y}(0,\cdot)$, $\widehat{\mu}^{A}(0,\cdot)$, $\widehat{\delta}$ are correctly specified, (ii) both $\widehat{\pi}$ and $\widehat{\delta}^{A}$ are, or (iii) $\widehat{\pi}$ and $\widehat{\delta}$ are. However, in the adaptive setting, we can establish a stronger result: even if the outcome-related nuisance functions $\widehat{\mu}_{t}^{Y}(0,\cdot)$ and $\widehat{\mu}_{t}^{A}(0,\cdot)$ are misspecified, AMRIV is still consistent as long as one of $\widehat{\delta}^{A}_{t}$ or $\widehat{\delta}_{t}$ converges. This is due to the adaptive setting design where we control the instrument assignment $\pi_{t}(X_{t}\mid\mathcal{H}_{t-1})$ which confers robustness to misspecification in $\widehat{\mu}_{t}^{Y}(0,\cdot)$ and $\widehat{\mu}_{t}^{A}(0,\cdot)$. Thus, our adaptive generalization preserves the multiple robustness property, making it particularly well-suited for practice where some nuisance components may be difficult to estimate reliably.
@@ 7 Experimental Results
Across both experiments, we evaluate three desiderata: (i) efficiency, measured by normalized MSE relative to the Oracle estimator; (ii) consistency, assessed via MSE decay with sample size $T$; and (iii) coverage, computed from empirical 95% confidence intervals. We implement all estimators using Random Forests (RF, [8]) and update the nuisance estimates in mini-batches for efficiency. Further implementation details, including model hyperparameters and ablations, are provided in Appendix F. The replication code is available at https://github.com/CausalML/Adaptive-IV.
@@ 7.1 Simulation Studies with Synthetic Data
We construct a synthetic environment with one-sided noncompliance, where the treatment $A$ is only accessible to those who receive the instrument $Z=1$. At each time $t$, we sample covariates $X_{t}$, assign the instrument $Z_{t}\sim\pi_{t}(X_{t}\mid\mathcal{H}_{t-1})$, and realize $A_{t}=C_{t}Z_{t}$, where $C_{t}$ is a latent compliance indicator sampled from $\text{Bern}(\delta^{A}(X_{t}))$. The outcome $Y_{t}$ depends on $A_{t}$, $X_{t}$, an unobserved confounder $U_{t}$, and heteroskedastic noise. The full data-generating process is detailed in Appendix F.1.
@@
We set $T=2000$, $T_{0}=200$, and run $1000$ trajectories. All estimators are updated in batches of size $b=200$ and implemented using Random Forests (RFs) when applicable. For the adaptive estimators, we use the truncated optimal policy in Eq. (7), with truncation schedule $k_{t}=2/0.999^{t}$. AMRIV uses RF classifiers for $\delta^{A}(X)$ (clipped at $0.01$) and RF regressors for $\mu^{Y}(z,X)$, while $\delta(X)$ is computed via plug-in. A2IPW follows Kato et al. [25] with Neyman allocation and RF regressors. To induce misspecification, we replace $\widehat{\mu}^{Y}(1,X)$ with the constant $\widehat{\mathbb{E}}[\mu^{Y}(1,X)]$, flattening outcome heterogeneity. Figure 2 summarizes the experimental results.
@@
Adaptivity. As shown in panel (a), adaptive design consistently improves the efficiency of all estimators. AMRIV approaches the oracle benchmark despite using estimated nuisances, while AMRIV-NA exhibits a constant efficiency gap due to suboptimal allocation. This illustrated the benefit of adaptivity in the one-sided noncompliance design, where asymmetries in outcome and compliance variance make uniform allocation especially inefficient (Theorem 1). Panel (a) also confirms that AMRIV and AMRIV-NA converge at the expected $O_{p}(T^{-1/2})$ rate (Theorem 4), whereas DM and DM-NA converge more slowly, as their normalized MSE increases with $T$.
@@
Consistency. Panel (b) confirms that AMRIV, AMRIV-NA, and DM converge to the true $\tau$, with AMRIV variants achieving lower error due to variance-aware allocation. In contrast, A2IPW is biased and fails to converge, as expected, since it does not correct for unobserved confounding in treatment selection. We also note the effect of misspecification: DM-MS does not converge due to incorrect estimation of $\delta(X)$, while AMRIV-MS remains consistent—highlighting the multiply-robust property formalized in Theorem 4, which allows consistency even when some nuisances are misspecified.
@@
Coverage. In panel (c), we evaluate the empirical coverage of 95% asymptotic confidence intervals. Only AMRIV and AMRIV-NA attain nominal coverage, consistent with our theoretical results (Theorem 3). Misspecified and plug-in methods under-cover, with DM and A2IPW performing particularly poorly with increasing $T$ due to finite-sample bias. AMRIV-MS provides partial correction but fails to reach nominal levels, as required by consistency of $\delta(X)$.
@@ 7.2 Simulation Studies with Semi-Synthetic Data
We also evaluate AMRIV on a semi-synthetic dataset based on the TripAdvisor customer simulator from Syrgkanis et al. [42], where we use customer features as covariates $X$, a simulated signup prompt as the instrument $Z$, and subscription revenue as the outcome $Y$. The DGP and oracle nuisances are described in Appendix F.2. Results are consistent with the synthetic setting: adaptive instrument assignment improves efficiency, AMRIV achieves superior coverage and consistency, and robustness holds under partial misspecification.
@@ 8 Conclusion
We develop AMRIV, an adaptive, multiply robust estimator for ATE estimation in experiments where treatment can only be encouraged via a binary instrument. Our approach (i) derives the semiparametric efficiency bound and optimal assignment policy, (ii) constructs a sequential estimator that attains the bound under adaptive allocation, (iii) provides asymptotic normality, convergence rates, and multiply robust consistency, and (iv) supports valid inference through time-uniform confidence sequences. Empirical results on synthetic and semi-synthetic data confirm that adaptive instrument assignment improves both efficiency and robustness over non-adaptive baselines. This work as a step toward principled, data-efficient experimentation in real-world settings where compliance is optional and uncertainty is unavoidable. We discuss limitations and broader impacts in Appendix G.
@@ A.1 Core Related Work
We instead build on the framework of Wang and Tchetgen Tchetgen [44], who introduce an alternative identification strategy based on an unconfounded compliance assumption. This allows point identification of the ATE while separating identification assumptions from estimation model assumptions. Their approach avoids reliance on parametric SEMs for ATE estimation and instead yields an efficient semiparametric estimator with an efficient influence function (EIF) that confers the estimator a multiple-robust property, i.e. the estimator remains consistent if one or several nuisance components are misspecified. This structure makes their estimators well-suited to nonparametric plug-in estimation using modern machine learning tools. Recent work extends this framework to nonparametric identification of ATEs under related assumptions [29].
@@
Adaptive Experimental Design for Treatment Effect Estimation. A growing body of work investigates adaptive algorithms for estimating the average treatment effect (ATE) efficiently and with minimal variance. This line of research was initiated by Hahn et al. [22], who proposed a two-stage design that asymptotically achieves the semiparametric efficiency bound, echoing explore-then-commit strategies in bandit settings. Building on this, Kato et al. [25] introduced the Adaptive Augmented Inverse Propensity Weighting (A2IPW) estimator, which enables fully sequential, covariate-dependent treatment assignment. They showed that this procedure achieves asymptotic optimality and outperforms earlier two-stage designs in finite-sample experiments.
Subsequent work has strengthened the theoretical and practical foundations of adaptive ATE estimation. Kato et al. [26] extended A2IPW to the setting where the assignment policy itself is estimated and demonstrated multiply robust consistency under partial nuisance misspecification. Cook et al. [13] further advanced this line by introducing principled policy truncation to control variance and ensure stable finite-sample behavior. They also established asymptotic normality and developed the first time-uniform confidence sequences for adaptive ATE estimation. Together, these contributions form the basis of a modern toolkit for adaptive experimentation that supports robust estimation, sequential inference, and strong theoretical guarantees.
Our work builds directly on this literature. We borrow and integrate many of its key components—influence-function–based estimation, cross-fitting, and policy truncation—to construct a complete, theoretically grounded estimator for experiments with noncompliance. Unlike existing work, however, we consider settings with endogenous treatment selection and unobserved confounding, where the experimenter can only assign a binary instrument rather than the treatment itself. To our knowledge, we provide the first adaptive estimator in this setting that achieves the semiparametric efficiency bound and supports valid sequential inference.
@@
The first group focuses on improving estimation efficiency in indicrect experiments. Gupta et al. [21] propose an adaptive framework for selecting among multiple data sources to efficiently estimate causal functionals such as the ATE. Their method, Online Moment Selection (OMS), chooses which source to query at each step based on moment conditions implied by a causal graph. While they address efficient data acquisition under structural constraints, their setting assumes passive data collection and differs from our focus on adaptive experimental design with noncompliance and endogenous treatment. Ailer et al. [3] study sequential indirect experiment design in instrumental variable settings, focusing on partial identification of nonlinear treatment effect queries. Rather than aiming for point estimation, their method adaptively tightens upper and lower bounds on a functional $Q[f]$ of the treatment effect by selecting experiments that reduce the gap between these bounds. In contrast, our work targets point identification and estimation of the ATE, and provides semiparametric efficiency and robustness guarantees under adaptive instrument assignment. Most closely related to our setting, Chandak et al. [9] study adaptive instrument selection to improve sample efficiency in indirect experiments. They propose a general influence-function–based optimization procedure for selecting instruments that minimize the mean squared error of nonparametric IV estimators, such as DeepIV [23]. However, their objective is variance reduction for prediction, not inference for causal estimands like the ATE. Their analysis is estimator-specific and does not characterize semiparametric efficiency bounds or multiply robust inference, which are central to our approach.

### v2

@@ 6 Theoretical Guarantees
This section provides theoretical guarantees for the AMRIV estimator. We establish its asymptotic normality, characterize its convergence rates, and demonstrate its multiply-robust consistency. Furthermore, in Appendix D, we consider the sequential inference setting and derive asymptotically-valid, time-uniform confidence sequences for the AMRIV estimator.
@@ 6.1 Efficiency and Asymptotic Normality of the AMRIV Estimator
Suppose Assumptions˜1, 2 and 3 hold and there exists a non-adaptive policy $\pi(X)\in[\epsilon,1-\epsilon]$ for some $\epsilon>0$ such that the nuisances estimates $\widehat{\eta}_{t}$ and the adaptive assignment policy $\pi_{t}(X\mid\mathcal{H}_{t-1})$ are $L_{2}$-consistent relative to the truncation schedule, i.e. $k_{t}\|\widehat{f}_{t-1}-f\|_{2}=o_{p}(1)$ and $k_{t}\|\pi_{t}-\pi\|_{2}=o_{p}(1)$ for $f\in\{\mu^{Y}(0,\cdot),\mu^{A}(0,\cdot),\delta(\cdot),\delta^{A}(\cdot)\}$. Furthemore, assume $\|\widehat{\delta}_{t-1}-\delta\|_{2}\|\widehat{\delta}^{A}_{t-1}-\delta^{A}\|_{2}=o_{p}(t^{-1/2})$. Then, the AMRIV estimator is asymptotically normal:
$\sqrt{T}\left(\widehat{\tau}^{\text{AMRIV}}_{T}-\tau\right)\overset{d}{\longrightarrow}\mathcal{N}\left(0,\,V_{\text{eff}}(\pi)\right),$ (10)
@@
The key insight behind Theorem 3 is that the AMRIV estimator admits the following near-martingale decomposition: $\sqrt{T}(\widehat{\tau}^{\text{AMRIV}}_{T}-\tau)=\sqrt{T}\left(\frac{1}{T}\sum_{t=1}^{T}z_{t}-\tau\right)+\sqrt{T}\left(\frac{1}{T}\sum_{t=1}^{T}m_{t}\right),$ where $z_{t}=\phi(X_{t},Z_{t},A_{t},Y_{t};\pi_{t},\eta)$ is a martingale difference sequence (MDS), and $m_{t}=\phi(X_{t},Z_{t},A_{t},Y_{t};\pi_{t},\widehat{\eta}_{t})-\phi(X_{t},Z_{t},A_{t},Y_{t};\pi_{t},\eta)$ is an asymptotically vanishing term that captures the impact of estimating the nuisance functions. The first term satisfies a central limit theorem for MDS under standard Lindeberg-type conditions [60], while the second is controlled by the $L_{2}$-consistency of the nuisance estimates. We formalize this in Appendix F. Importantly, this result holds under mild assumptions: we only require $L_{2}$ convergence (no pointwise convergence or Donsker conditions [7]), bounded outcomes and nuisance estimates, and $L_{2}$-consistency of the nuisance components w.r.t. the truncation schedule. This allows AMRIV to accommodate flexible, data-dependent policies and sequential nuisance estimation.
The truncation schedule $k_{t}$ plays a central role by ensuring positivity of $\pi_{t}(X)$—crucial for variance control—while still allowing $\pi_{t}$ to approach an optimal policy $\pi^{*}$ as $k_{t}\to\infty$. For this to hold, we must ensure $\lim_{t\to\infty}k_{t}>\sup_{X}\frac{1}{\pi^{*}(X)}$, so the truncation threshold does not constrain the optimal allocation in the limit and semiparametric efficiency can be achieved (see last line in Theorem 3). This mirrors tradeoffs in efficient ATE estimation [13], where adaptive truncation stabilizes estimation without distorting the estimator asymptotically. In practice, when the plug-in policy $\tilde{\pi}_{t}$ is uniformly bounded away from 0 and $1$, truncation becomes unnecessary: setting $k_{t}=1/\min_{X}\{\tilde{\pi}_{t}(X),1-\tilde{\pi}_{t}(X)\}$ ensures $\pi_{t}=\tilde{\pi}_{t}$ for all $t$.
@@ 6.2 Consistency Guarantees under Partial Nuisance Misspecification
Suppose Assumptions˜1, 2 and 3 hold, and that there exists a non-adaptive policy $\pi(X)\in[\epsilon,1-\epsilon]$ for some $\epsilon>0$ such that the adaptive policies $\pi_{t}$ satisfy $k_{t}\|\pi_{t}-\pi\|_{2}=o_{p}(1)$. Let $\widetilde{\eta}=\{\widetilde{\mu}^{Y}(0,\cdot),\widetilde{\mu}^{A}(0,\cdot),\widetilde{\delta}(\cdot),\widetilde{\delta}^{A}(\cdot)\}$ denote a possibly misspecified limit of the nuisance functions, and suppose that $k_{t}\|\widehat{f}_{t-1}-\widetilde{f}\|_{2}=o_{p}(1)$ for each $\widetilde{f}\in\widetilde{\eta}$. Then the AMRIV estimator satisfies
@@
We provide a proof of Theorem 4 and Corollary 5 in Appendix G. An immediate consequence of Theorem 4 is that if both $\widehat{\delta}(X)$ and $\widehat{\delta}^{A}(X)$ converge at rate $o_{p}(T^{-1/4})$, AMRIV achieves the parametric $O_{p}(T^{-1/2})$ rate. This is usually attainable under mild regularity conditions, even with flexible nonparametric models. Furthermore, Corollary 5 shows that AMRIV inherits the multiply robust property from its static counterpart [54, 18]. In the static setting, the MRIV converges if either (i) $\widehat{\mu}^{Y}(0,\cdot)$, $\widehat{\mu}^{A}(0,\cdot)$, $\widehat{\delta}$ are correctly specified, (ii) both $\widehat{\pi}$ and $\widehat{\delta}^{A}$ are, or (iii) $\widehat{\pi}$ and $\widehat{\delta}$ are. However, in the adaptive setting, we can establish a stronger result: even if the outcome-related nuisance functions $\widehat{\mu}_{t}^{Y}(0,\cdot)$ and $\widehat{\mu}_{t}^{A}(0,\cdot)$ are misspecified, AMRIV is still consistent as long as one of $\widehat{\delta}^{A}_{t}$ or $\widehat{\delta}_{t}$ converges. This is due to the adaptive setting design where we control the instrument assignment $\pi_{t}(X_{t}\mid\mathcal{H}_{t-1})$ which confers robustness to misspecification in $\widehat{\mu}_{t}^{Y}(0,\cdot)$ and $\widehat{\mu}_{t}^{A}(0,\cdot)$. Thus, our adaptive generalization preserves the multiple robustness property, making it particularly well-suited for practice where some nuisance components may be difficult to estimate reliably.
@@ 7 Experimental Results
Across both experiments, we evaluate three desiderata: (i) efficiency, measured by normalized MSE relative to the Oracle estimator; (ii) consistency, assessed via MSE decay with sample size $T$; and (iii) coverage, computed from empirical 95% confidence intervals. We implement all estimators using Random Forests (RF, [8]) and update the nuisance estimates in mini-batches for efficiency. Further implementation details, including model hyperparameters and ablations, are provided in Appendix H. The replication code is available at https://github.com/CausalML/Adaptive-IV.
@@ 7.1 Simulation Studies with Synthetic Data
We construct a synthetic environment with one-sided noncompliance, where the treatment $A$ is only accessible to those who receive the instrument $Z=1$. At each time $t$, we sample covariates $X_{t}$, assign the instrument $Z_{t}\sim\pi_{t}(X_{t}\mid\mathcal{H}_{t-1})$, and realize $A_{t}=C_{t}Z_{t}$, where $C_{t}$ is a latent compliance indicator sampled from $\text{Bern}(\delta^{A}(X_{t}))$. The outcome $Y_{t}$ depends on $A_{t}$, $X_{t}$, an unobserved confounder $U_{t}$, and heteroskedastic noise. The full data-generating process is detailed in Appendix H.1.
@@
We set $T=2000$, $T_{0}=200$, and run $1000$ trajectories. All estimators are updated in batches of size $b=200$ and implemented using Random Forests (RFs) when applicable. For the adaptive estimators, we use the truncated optimal policy in Eq. (7), with truncation schedule $k_{t}=2/0.999^{t}$. AMRIV uses RF classifiers for $\delta^{A}(X)$ (clipped at $0.01$) and RF regressors for $\mu^{Y}(z,X)$, while $\delta(X)$ is computed via the plug-in ratio. A2IPW follows Kato et al. [27] with Neyman allocation and RF regressors. To induce misspecification, we replace $\widehat{\mu}^{Y}(1,X)$ with the constant $\widehat{\mathbb{E}}[\mu^{Y}(1,X)]$, flattening outcome heterogeneity. Figure 3 summarizes the experimental results.
@@
Adaptivity. As shown in panel (a), adaptive design consistently improves the efficiency of all estimators. AMRIV approaches the oracle benchmark despite using estimated nuisances, while AMRIV-NA exhibits a constant efficiency gap due to suboptimal allocation. This illustrates how adaptivity enables more effective data collection: by dynamically allocating instruments to regions of high uncertainty, AMRIV concentrates sampling effort where it contributes most to precision. The effect is particularly evident under one-sided noncompliance, where asymmetries in both outcome and compliance variance make uniform allocation especially inefficient (Theorem 1). Panel (a) also confirms that AMRIV and AMRIV-NA converge at the expected $O_{p}(T^{-1/2})$ rate (Theorem 4), whereas DM and DM-NA converge more slowly, as their normalized MSE increases with $T$.
@@
Consistency. Panel (b) confirms that AMRIV, AMRIV-NA, and DM converge to the true $\tau$, with AMRIV variants achieving lower error due to variance-aware allocation. In contrast, A2IPW is biased and fails to converge, as expected, since it does not correct for unobserved confounding in treatment selection. The comparison further highlights the importance of robust nuisance estimation: DM-MS diverges when $\delta(X)$ is misspecified, while AMRIV-MS remains consistent. This robustness reflects the multiply-robust property formalized in Theorem 4, which ensures consistency as long as either the compliance model or the outcome model is estimated correctly—an especially desirable feature in practice when some nuisance components are difficult to learn reliably.
@@
Coverage. In panel (c), we evaluate the empirical coverage of 95% asymptotic confidence intervals. Only AMRIV and AMRIV-NA achieve nominal coverage, consistent with our theoretical guarantees (Theorem 3). The misspecified and plug-in methods under-cover, with DM and A2IPW performing particularly poorly as $T$ grows, owing to finite-sample bias and unaddressed confounding bias, respectively. AMRIV-MS provides partial correction but still falls short of nominal coverage, consistent with the requirement that $\delta(X)$ be estimated consistently for asymptotic validity.
@@ 7.2 Simulation Studies with Semi-Synthetic Data
We also evaluate AMRIV on a semi-synthetic dataset based on the TripAdvisor customer simulator from Syrgkanis et al. [52], where we use customer features as covariates $X$, a simulated signup prompt as the instrument $Z$, and subscription revenue as the outcome $Y$. The DGP and oracle nuisances are described in Appendix H.2. Results are consistent with the synthetic setting: adaptive instrument assignment improves efficiency, AMRIV achieves superior coverage and consistency, and robustness holds under partial misspecification.
Overall, these findings confirm that pairing adaptive design with the AMRIV estimator improves efficiency, enhances robustness, and yields more reliable inference than non-adaptive baselines.
@@ 8 Conclusion
We develop AMRIV, an adaptive, multiply robust estimator for ATE estimation in experiments where treatment can only be encouraged via a binary instrument. Our approach (i) derives the semiparametric efficiency bound and optimal assignment policy, (ii) constructs a sequential estimator that attains the bound under adaptive allocation, (iii) provides asymptotic normality, convergence rates, and multiply robust consistency, and (iv) supports valid inference through time-uniform confidence sequences. Empirical results on synthetic and semi-synthetic data confirm that adaptive instrument assignment improves both efficiency and robustness over non-adaptive baselines. This work represents a step toward principled, data-efficient experimentation in real-world settings where compliance is optional and uncertainty is unavoidable. We discuss limitations and broader impacts in Appendix I.
Acknowledgements
We thank the anonymous reviewers for their thoughtful feedback and insightful suggestions, which have helped improve and strengthen this work. Miruna Oprescu was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award DE-SC0023112. Brian M. Cho was supported by the Department of Defense (DoD) through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. Nathan Kallus was supported by the U.S. National Science Foundation under Grant No. 1846210. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. Department of Energy, the U.S. Department of Defense, or the U.S. National Science Foundation.
@@ A.1 Core Related Work
We instead build on the framework of Wang and Tchetgen Tchetgen [54], who introduce an alternative identification strategy based on an unconfounded compliance assumption. This allows point identification of the ATE while separating identification assumptions from estimation model assumptions. Their approach avoids reliance on parametric SEMs for ATE estimation and instead yields an efficient semiparametric estimator with an efficient influence function (EIF) that confers the estimator a multiple-robust property, i.e. the estimator remains consistent if one or several nuisance components are misspecified. This structure makes the estimators well-suited to nonparametric plug-in estimation using modern machine learning tools. This insight has been extended to develop multiply-robust CATE estimators that leverage binary instruments to adjust for unobserved confounding [18], and to debias confounded observational data by incorporating (potentially weak or imperfect) instruments [43]. Other recent work extends the framework of Wang and Tchetgen Tchetgen [54] to nonparametric identification of ATEs under related assumptions [34].
@@
Adaptive Experimental Design for Treatment Effect Estimation. A substantial literature studies adaptive algorithms for estimating the average treatment effect (ATE) efficiently and with minimal variance when the treatment itself can be directly assigned. This line of work was initiated by Hahn et al. [23], who proposed a two-stage explore-then-commit design that asymptotically achieves the semiparametric efficiency bound, echoing early bandit-style adaptive allocation schemes. Fully sequential designs soon followed: Kato et al. [27] introduced the Adaptive Augmented Inverse Propensity Weighting (A2IPW) estimator, which achieves variance-optimal Neyman allocation; Kato et al. [28] extended this to settings with estimated policies and nuisance functions and demonstrated multiply robust consistency; and Cook et al. [13] added principled policy truncation and developed the first anytime-valid confidence sequences for adaptive ATE estimation.
Parallel progress has come from an online-learning perspective. Recent methods such as Clip-OGD and its optimistic variants attain sublinear or logarithmic "Neyman regret" for the ATE [14, 38, 39]; low-switching policies achieve finite-sample optimality bounds [35]; and newer designs jointly optimize over covariate and treatment dimensions [29]. Off-policy estimators with adaptively collected data have also obtained sharp error rates and regret guarantees [33]. Together, these advances form a mature toolkit for adaptive experimentation that combines adaptive nuisance learning, cross-fitting, policy truncation, regret-minimizing allocation, and time-uniform inference.
Our work builds directly on this literature but extends it to the far less explored setting where only an instrument can be assigned and treatment uptake is endogenous. We integrate core components from the direct-treatment literature—influence-function–based estimation, adaptive policy learning, cross-fitting, policy truncation, and sequential inference—to construct a unified estimator that retains semiparametric efficiency, multiply robust consistency, and time-uniform inference under noncompliance. Specifically, we are the first to: (i) derive the semiparametric efficiency bound for ATE estimation when only an instrument can be adaptively assigned; (ii) identify the variance-optimal allocation rule that balances outcome noise and compliance variability; and (iii) develop AMRIV, a multiply robust estimator that attains this bound and supports anytime-valid inference.
@@
The first group focuses on improving estimation efficiency in indirect experiments. Gupta et al. [22] propose an adaptive framework for selecting among multiple data sources to efficiently estimate causal functionals such as the ATE. Their method, Online Moment Selection (OMS), chooses which source to query at each step based on moment conditions implied by a causal graph. While they address efficient data acquisition under structural constraints, their setting assumes passive data collection and differs from our focus on adaptive experimental design with noncompliance and endogenous treatment. Ailer et al. [3] study sequential indirect experiment design in instrumental variable settings, focusing on partial identification of nonlinear treatment effect queries. Rather than aiming for point estimation, their method adaptively tightens upper and lower bounds on a functional $Q[f]$ of the treatment effect by selecting experiments that reduce the gap between these bounds. In contrast, our work targets point identification and estimation of the ATE, and provides semiparametric efficiency and robustness guarantees under adaptive instrument assignment. Most closely related to our setting, Chandak et al. [9] study adaptive instrument selection to improve sample efficiency in indirect experiments. They propose a general influence-function–based optimization procedure for selecting instruments that minimize the mean squared error of nonparametric IV estimators, such as DeepIV [24]. However, their objective is variance reduction for prediction, not inference for causal estimands like the ATE. Their analysis is estimator-specific and does not characterize semiparametric efficiency bounds or multiply robust inference, which are central to our approach.